Cinema seats are numbered with those numbers: 1, 2, ..., 20. 6 friends one after another buys tickets and seats are allocated to them at random order.
a) What are the chances that everyone will get seats with even numbers?
b) 2 friends that come earlier got seats with even numbers. What are the probability that the rest will also get places with even numbers?

Answer by greenestamps(13203)



(a) At the start, there are 10 even-numbered seats out of a total of 20.

P(1st seat # is even) = 10/20
P(2nd seat # is even) = 9/19
...
P(5th seat # is even) = 6/16
P(6th seat # is even) = 5/15

You want all those things to happen, so multiply all those probabilities. I leave it to you to do the calculations and express the result in the required format.

(b) In this case, 2 of the 10 even-numbered seats are already occupied. The solution process is the same as in part (a); but now the probabilities are 8/18, 7/17, ..., 4/14, and 3/13.